OpenStudy (anonymous):

The table below shows two equations: Equation 1 |4x - 3|- 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations? (4 points) Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1.

OpenStudy (jdoe0001):

\(\large {|4x - 3|- 5 = 4\implies |4x - 3|=4+5\implies|4x - 3|=9\\ \quad \\ |4x - 3|=9\implies \begin{cases} +(4x - 3)=9\\ \quad \\ -(4x - 3)=9 \end{cases}\\ \quad ------------------------\\ |2x + 3| + 8 = 3\implies|2x + 3| = 3-8\implies|2x + 3| =-5\\ \quad \\ |2x + 3| =-5\implies \begin{cases} +(2x + 3) =-5\\ \quad \\ -(2x + 3) =-5 \end{cases}}\)